OpenStudy (anonymous):

solve this system of equation using the addition method? 4a - 5b = 7 4a + 5b = 17

OpenStudy (anonymous):

4a-5b=7 4a+5b=17 add these two equations and you get 8a=24 divide both sides by 8 and you get a=3

OpenStudy (anonymous):

oh okay thanks that help :) ^^

OpenStudy (anonymous):

what about when you have one like this one ?? x + y = 8 -x + 2y = 7

OpenStudy (anonymous):

again add both the equations and you will get 3y=15 divide both side by 3 and you get y=5 once you have this value so can work out the value of x in this example as well as the previous example

OpenStudy (anonymous):

thanks :) ^^